Night report 2017-11-27

Observers: Z.Kostrzewa-Rutkowska, K.Lopez

TAT: J. Chen, M. Fraser, J. Sollerman

Support team: P. Jonker, G. Cannizzaro, F. Onori

Instruments: SOFI, EFOSC2

Weather conditions:

evening UT 23:45: clear, humidity 29%

strong wind

morning UT 8:30: thin clouds, humidity 25%

Calibrations:

EFOSC2 bias, arcs daytime, arcs morning

LOG:

UT Target Type Filter/Grism ExpTime[s] Slit Seeing Airmass Comments

--------SOFI---------

00:03 LTT7987 STD BGRG ~0.8 1.42

00:18 AT2016iks Followup JHK ~0.8 1.23

01:30 Technical issues

01:50 MASTERJ034412.. Followup JHK 1.11 1.13

03:12 P9101 STD JHK ~1.5 1.64

03:17 HIP10545 Telluric BGRG 1.62 1.05

03:23 AT2017gmr Followup BGRG 1.61 1.07

04:40 SOFI --> EFOSC2

-------EFOSC2--------

04:50 L745a STD gr11/gr13/gr16 1.0/1.5 1.40 1.38

05:20 AT2017ilt Class gr13 600 1.0 1.26 1.88

05:38 AT2017ilv Class gr13 600 1.0 1.40 1.83

05:55 AT2017iit Followup gr11 1800 1.0 1.36 1.52

06:35 AT2017hpi Followup gr13 1500 1.0 ~1.8 1.08 Fixing focus

06:48 AT2017hpi Followup gr13 1500 1.0 1.20 1.07 Restarted

07:20 AT2017ile Class gr13 300 1.0 1.60 1.07

07:34 AT2017imd Class gr13 600 1.5 1.93 1.39

07:53 AT2017ijq Class gr13 900 1.0 1.30

08:17 L745a STD gr11/gr13/gr16 1.0/1.5 0.93 1.02